which of these points lies on the line described by the equation below? y-8=3(x+6)

Answer by Alan3354(69443)

which of these points lies on the line described by the equation below? y-8=3(x+6)
================
No points are shown.
y-8=3(x+6)
y = 3x + 26
------
Sub for x & y of the points, see if they fit.